### v1

Abstract
Artificial engineering often relies on scarce materials, possessing ecological, social, and political risks. In contrast, nature achieves advanced functionalities using only abundant elements and water. Despite this simple insight, incorporating nature-inspired devices made from readily available resources into industrial applications remains a formidable challenge. Here, we present a sustainable ’blue battery’—a water-based energy storage device that mimics natural processes and reduces dependence on scarce materials. This battery harnesses the distinctive electrical properties of water confined within nanometer-scale pores, enabling construction from abundant, non-toxic materials using scalable nanofabrication techniques. It operates without unwanted side reactions, maintains nearly 100% efficiency after 60,000 charge-discharge cycles, and supports a voltage window of up to 1.65 V while achieving competitive power and energy density. The core component, a van der Waals clay nanostructure, uses water as the sole electrolyte with enhanced activity due to interactions with its 2D surface. This water-driven technology offers a sustainable platform for energy systems, with potential applications from small-scale electronics to global energy infrastructure—even in extreme environments like Mars. Our work advances nature-inspired energy solutions, opening new possibilities for the sustainable engineering of blue devices.
1 Main
In contrast to man-made engineering approaches, nature builds its functionalities using abundant ’primitive’ chemical elements and water, prioritizing nanoscale structuring over the optimization of chemical composition [1, 2, 3]. This strategy, which ensures scalability and growth critical for survival [4], presents a blueprint for engineering innovations aimed at sustainability. By mimicking nature, we can reduce dependence on scarce materials and develop water-based systems—hereafter referred to as ’blue devices’—which hold the potential to meet rising demands in areas such as sustainable batteries [5], neurointerfaces [6], and neuromorphic computing [7]. However, a thorough understanding of water’s atomic and molecular behavior at interfaces, a topic of growing interest in recent years [8, 9, 10, 11, 12, 13], remains essential for advancing these technologies.
Recent studies have revealed unusual properties of water confined in pores ranging from a few angstroms to several nanometers, or ultraconfinement [14, 15, 16]. Under these extreme conditions, water exhibits behavior distinct from bulk water, challenging existing theoretical models [17]. Experiments have revealed unique effects such as dielectric polarization anisotropy and protonic ’superconductivity,’ both linked to altered proton dynamics in ultraconfined water [18, 19, 20, 21]. These discoveries open exciting possibilities for blue devices, particularly in improving aqueous batteries and supercapacitors. These devices offer safer, more cost-effective alternatives to traditional technologies that rely on metal oxides and concentrated electrolytes [22, 23], and are crucial for stabilizing renewable energy networks powered by solar and wind. While previous efforts in nanofluidic systems have demonstrated impressive performance in energy applications [24, 25], scaling them up for practical use remains a significant challenge [26, 27]. Utilizing ultraconfined water in nature-inspired designs with abundant materials could unlock new scaling-up possibilities while addressing environmental and sustainability concerns.
In this work, inspired by nature, we introduce a sustainable ’blue battery.’ This water-based device utilizes pure water confined within extremely narrow 1-nm-high slit pores as an electrolyte, where proton activity is significantly enhanced compared to bulk water. We demonstrate that few-molecule-high water channels facilitate the separation and accumulation of protons and hydroxides within the electrodes’ electrical double layers (EDLs). To ensure practical applicability, the device is constructed exclusively from abundant materials—clay, graphene, and water—reducing reliance on scarce resources in energy storage. The device achieves up to 100% coulombic efficiency, with a projected equivalent lifespan of 150 years without significant degradation. It operates within a voltage window of 1.65 V, offering competitive power and energy densities. By focusing on precise structural design rather than rare chemical elements, our ’blue battery’ provides a scalable and sustainable platform for various future technologies. These include safe energy storage systems, electrolyzers, bio-computer interfaces, and neuromorphic devices, where ultraconfined water within solids serves as both an electrolyte and a separator.
1.1 Topologically connected channels as blue device concept
Unlike previous aqueous electrochemical systems with geometrically disjointed electrodes (Fig. 1, a and b), we introduce a topologically connected membrane-electrode unit (Fig. 1c), which mimics natural systems such as the interstitial fluid in biological organisms (Fig. 1d), providing structural continuity within tissue and organ boundaries [28]. Consider the simplest aqueous capacitor, the ’Leyden jar’ [29], which consists of two metal electrodes in contact with water (Fig. 1a). In this device, electric energy is stored in a thin, nanometer-scale interface (blue glow) with a low energy density due to a negligible surface-to-volume ratio. To improve this, the classic battery or supercapacitor design (Fig. 1b) increases the surface area of the electrodes to enhance energy density but also increases an electrolyte concentration to secure high current density, thereby promoting chemical complexity. In contrast, we propose transforming the disjointed electrodes into a unified, topologically connected network of electrical double layers (EDLs) without involving bulk water (Fig. 1c). This design leverages the unique properties of ultraconfined water without increasing chemical complexity, maintaining sustainability without compromising performance.
To design our blue device, we created an electrode-membrane-electrode unit out of abundant 2D materials loaded with ultraconfined water within 1-nm-high slit channels (Fig. 1e). In the channels the walls are so close that the double layers of water from opposite sides overlap, eliminating bulk water and leaving only interfacial water with affected self-dissociation-recombination balance [17]. Such water exhibits modified electrodynamic properties and an increased activity of H+ and OH- ions. Additionally, the EDLs of the electrodes connect topologically through the EDLs of the separator, creating a continuous network across the device parts. This design mimics nature by focusing on space structuring rather than chemical diversity. As demonstrated below, the device achieves exceptional cycle life and competitive performance. Water access into the aligned slit pores (Fig. 1f) is ensured by the use of hydrophilic clay, whose properties require a separate analysis at the microscopic level.
1.2 Where water meets Earth: clays as natural 2D materials
Clays occur on all Earth’s continents (Fig. 2a) and also on Mars [30]. On the microscopic level, they have fine crystalline structure (Fig. 2b) [31]. The most common types of clay—kaolinite, illite, and smectite (Fig. 2c)—consist of 1-nm-thick crystals made of silica and alumina planes, interchanging differently depending on the clay type to form flat flakes. These flakes are held together by vdW interactions, forming periodic heterostructures up to micrometers in size (SI Figs. S1-S4). Natural clays are chemically simple and composed of silicon, oxygen, and aluminum atoms but may also include water-soluble potassium, sodium, or calcium, which can be reversibly substituted by, e.g. hydrogen [31, 32].
Unlike synthetic vdW heterostructures [34, 35], clays offer several advantages: low price, scalability, widespread availability, natural affinity to water, safety, and biocompatibility. Their organized slit channels (Fig. 2b) facilitate water and proton transport [32], allowing for the movement and relative separation of positive and negative ions under natural conditions or an external electric field. Besides intercellular spaces (Fig. 1d), the clay interlayer pores are similar to small wood pores or nerve cell structures [36, 37], making them suitable for nature-mimicking engineering.
Our X-ray diffraction (XRD) analysis showed that exposing dried clays to saturated water vapor or liquid water results in interlayer expansion or swelling, evidenced by diffraction peak shifts (Fig. 2e and SI Figs. S5 and S6). This process and the natural clay structure allow the creation of periodic slit channels filled with water (Fig. 2d). Notably, this procedure is much less expensive and time-consuming than the artificial vdW channels design [38] and has almost infinite materials sourcing capabilities. We found that smectite, in particular, has an equilibrium pore height of 1 nm, consistent with previous reports [39, 40]. Coincidentally, these dimensions correspond to the peak maximum of proton conductivity observed in water confined in a nanoporous medium and within artificial slit pores [18, 20, 21].
To test the electrical properties of clays, we first carefully cleaned the samples and then hydrated them in saturated water vapor (see Methods for sample preparation). This process ensured the absence of contamination and foreign ionic impurities. Dry clays acted as good insulators, while wet clays conducted ions with the assistance of water. We observed significant ionic conductivity in all wet clays, in contrast to both bulk water and dry clays, which exhibited high resistivity (1 Mcm, Fig. 2f). Smectite, in particular, demonstrated the highest direct-current (DC) conductivity of 0.6 S/m, comparable to that found in synthetic slits and ceramics [21, 20], and equivalent to the conductivity of a 0.1-M NaCl solution in bulk water or ionic liquids [41]. Notably, this conductivity is six orders of magnitude greater than that of bulk water, which is 0.610-7 S/m [17] (see blue dashed line).
The enhanced ionic transport in hydrated clay can be understood via dynamic interactions between the water and the clay under conditions of ultraconfinement. The natural substitution of Si4+ and Al3+ atoms in the clay’s crystal structure by Al3+ and Mg2+, respectively, creates an excess negative surface charge [31]. In nature, this excess surface charge is balanced by adsorbed alkaline cations to maintain electro-neutrality [31]. In cleaned clays, this role is played by H+ (excess protons) of water, whose adsorption and desorption increase the conductivity of confined water by disrupting its self-dissociation equilibrium [42] and raising the concentration and mobility of H3O+ and OH- (see inset in Fig. 2f). This effect is most pronounced in pores around 1.5 nm (approximately 2 times the Bjerrum length, nm), aligning with experimental data [20, 21] and predictions from the ionic model of water [17]. Notably, excess protons move via a relay-race proton-hopping mechanism, with water molecules acting as mediators, resulting in significantly higher mobility compared to other ions [17].
This scenario is supported by the negative surface charge we detected on colloidal nanoparticles of clays in distilled water (SI Table S1). The interaction between clay and water enables ultrahigh ion conduction within the 1-nm slit channels, transforming confined water into a partially ionic liquid with H+ and OH- ions as charge carriers. This effect, observed in other nanoconfinement systems [18, 20, 21], provides opportunities for nature-inspired engineering of blue devices. The use of clay, which processing does not require any extra chemicals unlike other 2D materials, ensures low cost and natural hydration of the system.
1.3 Energy storage in confined water: blue battery
To utilize the unique properties of ultraconfined water, including its inherent safety, giant in-plane polarizability [21], and enhanced conductance from surface-water interactions [20], we designed an electricity storage unit (blue battery). Using the nanopore vacuum filtration technique (Fig. 3a; see Methods), we sequentially and controllably self-assembled layer by layer the mixed graphene-clay vdW heterostructures to form two composite electrodes divided by a pure clay separator. The result was a scalable, free-standing membrane-electrode unit (MEU) in the form of nanocomposite film approximately 200 m thick (Fig. 3b, c), featuring a three-layer structure (Fig. 3d), and having high stability and flexibility, typical for vdW heterostructures [38].
SEM EDX mapping (Fig. 3e, f) confirmed the presence of only clay (O, Al, Si) and graphene (C) elements (see SI Figs. S9, S10 for additional data). Electron microscopy showed that within the layers (Fig. 3g, h), material flakes are aligned to form parallel channels, which intersect at the edges of the domains (Fig. 3i). Moreover, no voids are detected at the interface between the domains of the separator and the electrodes. The channels were filled with water via capillary condensation from saturated water vapor, with mass gain used to control the process. This design promotes ultraconfined water intercalation into the electrodes and the separator while eliminating bulk-water voids. The structure aligns with the topological design shown in Fig. 1c. Notably, this scalable system requires no additional chemicals and is highly reproducible using standard laboratory procedures.
To test the device, two graphite current collectors were attached to both sides of the MEU (Fig. 4a). The cell underwent standard electrochemical tests within a 2.1 V voltage window (see Methods) to evaluate charge-discharge cycles (Fig. 4b), current-voltage characteristics (Fig. 4c), and estimate capacitance, coulombic efficiency, and energy efficiency. Despite its simple design (Fig. 4d) and use of pure water without additives, the cell showed impressive parameters and durability with no visible degradation after 60,000 cycles (Fig. 4e). This cycle life translates to over 150 years of operation counting one charge-discharge cycle per day, e.g., in the case of use with solar panels. It retains and releases energy with up to 100% coulombic and 60% energy efficiency, due to the absence of side reactions. The initial ”annealing” period was assigned to the electric-field-stimulated wetting of nanopores.
Moreover, we found that the device has an electrolysis threshold of 1.65 V (Fig. 4f, and Fig. S11), higher than the standard 1.23-V limit for bulk water at neutral conditions, which yields higher capacitance, and energy density than that in bulk water systems. This is likely due to the combined effects of high water purity and the ultraconfinement of water, which alters its electrodynamics, particularly conductivity (Fig. 2f) and polarizability [18, 19, 21]. Our system achieves a capacitance of around 40 F/g and an energy density of around 10 Wh/kg of electrode material, comparable to that of commercial supercapacitors [43]. Both parameters have the potential for further improvement through optimization.
Note that the device’s efficiency and energy density significantly exceed those of biological systems such as photosynthesis, which has an efficiency below 1% [44]. Thus, our device mimics, but also surpasses nature. It traps and transports protons and hydroxides (proton holes) of the nanometer-thick water-solid interface within the cell over macroscopic distances in an external electric field and stores them in the same-size nanopores in the electrodes (see Fig. 1e). This effect is universal and can be used in other types of electrochemical systems, such as electrolyzers, batteries, and fuel cells, as well as nature-signaling mimicking systems and neuromorphic devices.
2 Discussion
We have demonstrated a scalable platform for energy applications that leverages the anomalous properties of ultraconfined water: an enhanced proton conductivity [20] and heterogeneous polarizability [21]. The system is topologically non-equivalent to the standard electrochemical systems design (Fig. 1, a to c) and mimics nature’s use of ultraconfinement and space (Fig. 1d). Our device, made from only abundant materials: graphene, clay, and water confined within 1-nm slit pores, serves as a rechargeable electricity retention unit with a projected lifespan of over 150 years (Fig. 4e). This makes it ideal for supplementing solar or wind generators to avoid grid instability, which is an urgent problem of a fast-growing renewable energy network. The exceptional durability and infinite materials sourcing suggest it could be a reliable solution for balancing electricity generation-consumption time lags or providing fast-frequency control in the grid [45]. A performance comparison with current batteries and supercapacitors indicates that our system could offer superior long-term stability that suggests specific niche applications (SI Table S3).
Note that while using rare elements in energy devices can enhance reaction speed, energy density, efficiency, and operational ranges (voltage or temperature), our nature-mimicking approach suggests advantages in terms of cost, environmental safety, geopolitical stability, biocompatibility, scalability, structural stability, and durability. The non-toxic nature of clay [46] makes our platform also suitable for a range of bio-applications, including battery pills [47], bio-interfaces [6], neuromorphic computing [7], and even for use in Mars colonization, where clays and water are available on-site [48], whereas ultraconfined water does not freeze down to -70 ∘C [49].
In conclusion, our blue-battery platform represents a significant advance in applying nature-inspired principles to sustainable nanoengineering of energy systems. By combining 2D materials science with abundant materials and simple fabrication, we achieve high performance, safety, material independence, and sustainability. Our technology aims to drive further research and seek sustainable industrial applications, promoting a more efficient, environmentally friendly, and safer energy future.
3 Methods
3.1 Sample preparation
Several clay types: bentonite, montmorillonite, illite, kaolinite, and silt, were purchased from Merck and Nagra in the form of powder and grains. The samples were preliminary milled to the sub-micron sizes. Then, clay samples were washed out through a multi-step cleaning process involving repeated dissolution, natural sedimentation preselection, washing and centrifugation cycles, dialysis, and sonication, with zero chemicals used. The resulting product was a nonprecipitating colloidal solution of clay nanoparticles in distilled water with a concentration of about 0.1 mg/ml, a bulk electrical resistivity of approximately 18 M, and a pH of 7. An electrochemically exfoliated graphene solution in distilled water was purchased from XFNano and used as received. Milli-Q water, with an 18 M resistivity, served as the water source. Polymer-filter substrates with nanopores ranging from 20 to 100 nm were purchased from Merck and used without further modification.
3.2 Vacuum filtration
Clay films and devices were fabricated using vacuum filtration of colloidal graphene and pre-treated clay solutions through the nanopore polymer-filter-substrate vacuum filtration. Before use, all solutions were sonicated at 1200 Watts and 20-25 kHz and naturally thermally equilibrated with the lab temperature. For electrode fabrication, either a pure graphene solution or a clay-graphene suspension mixture was used (see SI Table S2 for different MEU compositions). A pure clay solution was utilized for the separator. To create the three-layer electrode-separator-electrode structure, the electrode solution, separator solution, and second electrode solution were sequentially filtered. A water vacuum pump facilitated water transport through the filter, and filtration proceeded step-by-step, waiting for the complete passage of water from the previous step. Conductivity and pH tests of the filtrate confirmed no contamination.
Unlike pressing dry powders, vacuum filtration produces highly dense samples without large pores, featuring unidirectionally oriented ”brick-and-mortar” structure layers of 2D materials held together by vdW forces. This method ensures remarkable mechanical stability, flexibility, and reproducibility for each membrane-electrode assembly.
3.3 Electrochemical measurements
Vacuum-filtrated samples were placed in a saturated water vapor atmosphere for 48 hours, with water saturation confirmed by mass gain and swelling checks, ensuring no contamination by foreign ionic or molecular species. Conductivity, charge-discharge (CD), cyclic voltammetry (CV), and long-term CD tests were conducted using the BioLogic SP-300 and CH Instruments devices. The voltage window was set to 2.1 V for preliminary tests and to 1.6 V for main measurements, including long-term stability tests. The second value was chosen according to the electrolysis limit observed via a hydrogen-evolution-reaction (HER) indicator in CV and CD curves (SI Fig. S11). Two graphite current collectors were applied to the electrodes in all measurements. To prevent water evaporation, Parafilm was used to seal the membrane-electrode unit. Notably, no mass change in the samples was detected over one month after the fabrication. Pressure applied to the current collectors was controlled by a custom-made sample holder and was adjusted to ensure good electrical contact.
3.4 X-ray diffraction measurement
X-ray diffraction (XRD) measurements were conducted using a Bruker D8 Discover in Bragg Brentano geometry, with a Kalpha1 - monochromated beam ( = 1.5406 ) and a Lynxeye-XE detector and at the P62 SAXS/WAXS beamline of PETRA III synchrotron facility at Deutsches Elektronen-Synchrotron (DESY). We prepared two types of samples for these measurements: clay vacuum-filtered membranes and powders. For the Bruker measurements, dry samples were prepared by vacuum annealing for 24 hours, while wet samples were prepared by exposure to saturated water vapor for 48 hours. At PETRA III, the dry samples were similarly annealed, but the wet samples were prepared by directly adding water to the powders. The crystalline interlayer distance change on water adding (swelling effect) was found in both the PETRA III and lab-based XRD measurements (SI Fig. S7).
3.5 Scanning electron microscopy
Scanning electron microscopy (SEM) observation including EDX (energy dispersive X-ray spectroscopy) measurements was performed on an SEM/FIB (focused ion beam) Zeiss CrossBeam 550 microscope equipped with Oxford Instruments Ultim Max detector. SEM images were taken with 5 kV (Figs. S21, S22) and 3 kV (Figs. S23) accelerating voltage, while EDX data measurements were done at 15 kV (Figs. 3, d to g, and Figs. S9, S10). For subsequent transmission electron microscopy observation, a lamella from a smectite membrane (Figs. S24, S25) was prepared using the SEM/FIB tool Zeiss NVision 40 using an accelerated Ga+ beam at 30 kV and 5 kV.
3.6 Scanning transmission electron microscopy
Scanning transmission electron microscopy (STEM) experiments were realized on an aberration-corrected FEI Titan Themis3 microscope operated at 300 kV. An electron probe of an 8 mrad half-convergence angle was used in association with a current of 10 pA. A dwell time of 2 µs was used for STEM imaging, which limited the total electron dose deposited on the sample for acquiring atomic-resolution STEM images to 460 e-/. Annular dark-field (ADF) and integrated differential phase contrast (iDPC) STEM images were collected simultaneously, using a collection angle ranging from 17 to 104 mrad for ADF and from 4 to 15 mrad for iDPC. iDPC STEM imaging allowed us to visualize the atomic structure of the crystal composed of light elements [50], with the moderate electron dose used to avoid degradation of the crystal structure during imaging.
4 Acknowledgements
We thank Keith Stevenson, Nianduo Cai, Artem Pronin, Maria Nieves López Salas, and Anton Andreev for fruitful discussions, Sylvio Haas (DESY) for help during the synchrotron radiation-based X-ray scattering experiments, Vytautas Navikas for the help with the representation of Figure 1e, and Milad Sabzehparvar for assistance in the lab. We also acknowledge the scientific exchange and support of the Centre for Molecular Water Science, and the research initiative BlueMat: Water-Driven Materials, Hamburg (Germany). We thank the EPFL Center for Electron Microscopy (CIME) for access to electron microscopes.
Funding
Y.T. acknowledges the financial support from the Swiss National Science Foundation through the National Centre of Competence in Research Bio-Inspired Materials and Grant No. 200021-192037. A.Ra. and T.H.C. acknowledge funding from the European Research Council (grant 101020445—2D-LIQUID). S.B. and L.L. are grateful for the financial support provided by the Swiss National Science Foundation (grant No. 200021-204099, Division II). Funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in the DFG Graduate School GRK 2462 ‘Processes in natural and technical Particle-Fluid-Systems (PintPFS)’ (Project No. 390794421) is gratefully acknowledged. G.T. acknowledges support from the SNSF Eccellenza Grant 194181 and the SNSF Starting Grant 211695
Declarations
There are no conflicts of interest to declare. All data is available within this paper and Supplementary Information.
Author contribution
V.A. proposed and directed the research with the help of F.B., A.Ra., G.T., and L.L.; S.B. prepared the clay materials and characterized their optical, transport, and electrical properties together with V.A. and under the supervision of L.L.; Y.T. developed the vacuum filtration protocol and assembled the membranes and membrane-electrode units with V.A.; J.M. and V.A. conducted the electrochemical measurement under the guidance of G.T.; A.Ry. carried out the data analysis; T.-H.C. develop the protocol and processed the raw materials, controlled their purity, and measured -potentials; L.N., V.B., and Y.T. conducted electron microscopy characterization; P.S. conducted X-ray diffraction experiments; M.L. and P.H. conducted synchrotron radiation-based X-ray scattering experiments and analyzed the data; V.A. wrote the manuscript; and all authors contributed to discussions.