### v3

Abstract
Nanometer-scale solid-state confinement has been shown to change water’s structure and dynamics, offering new horizons in energy storage. However, most current materials operate at the micrometer scale, missing the interfacial effects that occur at three orders of magnitude smaller dimensions. Here, we report a scalable energy storage device that uses ultraconfined water as its sole electrolyte, unlocking the advantages of nanoscale confinement. We use the polarizability and proton ’superconductivity’ of water confined in few-molecular-diameters clay channels to build an all-water supercapacitor. The device fabricated from reconstructed clay, graphene, and water by a sustainable self-assembly process, operates at voltages up to 1.65 V, has competitive power and energy density, and maintains near 100% Coulombic efficiency over 60,000 charge-discharge cycles. These results demonstrate the application of unique properties of ultraconfined water for sustainable energy storage and provide a benchmark for a class of novel ultraconfined water-based energy systems, or ’blue devices’.
1 Introduction
The ultraconfinement of liquids within solid-state nanopores has gained exceptional attention over the past decade due to its critical role in emerging technologies, including osmotic power generation, ultrafiltration, and molecular sensing [1, 2, 3, 4, 5]. A liquid becomes ultraconfined when the electrical double layers (EDLs) from opposing surfaces overlap. In these conditions, the liquid exhibits drastically different—and sometimes counterintuitive—electrodynamic properties compared to bulk liquid [6]. The deviations in the rheological, mechanical, and electrodynamic properties of ultraconfined liquids in the pores of a few nanometers are also evident from the myriads of natural interfacial phenomena [7, 8]. Despite the extensive fundamental studies [9, 10, 11], the electrodynamic properties of ultraconfined liquids remain essentially unexploited because the nanochannels fabrication and measurements at the nanoscale are challenging.
Recent studies have shown dielectric anomalies in ultraconfined pure water, revealing structural heterogeneity and unusual molecular dynamics. Specifically, the dielectric constant anisotropy and the effect of proton ’superconductivity’ of ultraconfined water have been experimentally observed [12, 13, 14, 15] and supported by simulations [16]. These effects challenge our understanding of water’s bulk properties and suggest intriguing applications. Notably, they open up possibilities for sustainable and safer electrochemical energy storage systems, such as supercapacitors, by utilizing the topology and strong confinement instead of increasing chemical complexity. This approach mirrors nature’s strategy of emphasizing shape and structure over increasing chemical species to develop functionality [17], and yet to be explored.
Unlike phenomena in traditional aqueous supercapacitors, which utilize water as a solvent for concentrated electrolytes, the ultraconfinement affects the intrinsic electrodynamic properties of water without changing its atomic composition. This phenomenon has been understood via the surface-water interaction, which affects its self-ionization followed by charge separation [18]. The altered electrodynamics of ultraconfined water has been observed in various nanoporous systems, including that of hydrophilic and hydrophobic properties [12, 14, 19, 15]. This means that the phenomenon has a universal nature on a broader range of functional materials beyond the polymers traditionally used to enhance proton transport [20]. Although an attempt was made to create an all-water electrochemical supercapacitor using nanodiamond ceramics [19], the limited scalability of this and other nanoporous platforms has restricted the application of ultraconfined water in large-scale electricity storage.
Here, we introduce a scalable energy storage platform, called the ’blue device’, which utilizes electrical double-layer capacitance of ultraconfined pure water. By harnessing the enhanced proton conductivity of up to 0.6 S/m within 1-nm water channels in nanoengineered naturally abundant clays, we achieve energy and power densities competitive with conventional supercapacitors, hybrid capacitors, and some aqueous batteries. Operating at 1.65 V with nearly 100% Coulombic efficiency and a projected lifespan of 150 years, the device offers a highly sustainable, safe, and eco-friendly solution for bulk electricity storage.
2 Results
2.1 Topologically connected channels and blue device concept
The simplest aqueous capacitor is the Leyden jar [21] (Fig. 1a). It consists of two non-porous electrodes in contact with water. The electric energy is stored in a thin nm-high interface (EDL). Though sustainable, the device has negligible for practical applications energy density due to a small surface-to-volume ratio and low concentration of charge carriers able to travel between the electrodes. In contrast, the modern aqueous supercapacitor has large-surface-area electrodes immersed in a concentrated electrolyte or an ionic liquid (Fig. 1b). It suggests a relatively high power and energy density yet has chemical complexity. The latter increases the price and raises sustainability issues.
To address these problems, we propose transforming the disjointed electrodes into a unified, topologically connected network of nanopores lasting continuously between and within the electrodes through the nanoporous separator without a break (Fig. 1c). This design leverages the unique properties of water ultraconfined in nanochannels. Particularly, benefiting from the proton ’superconductivity’ and anomalous polarizability without increasing chemical complexity. The design is topologically nonequivalent to the previous two and maintains sustainability like the Leyden jar, but without compromising performance compared to standard supercapacitors.
A corresponding to the sketch in Fig. 1c device membrane-electrode unit (MEU) (Fig. 2a) can be assembled by using abundant clay and graphene nanosheets. This design suggests a continuum of water-filled nanochannels (Fig. 2b) between periodic solid-state sheets.
Unlike synthetic vdW heterostructures [22, 23, 24], natural clay nanosheets offer several advantages. They are low-cost, scalable, widely available, water-affine, safe, and biocompatible. Slit-like channels between the clay sheets (Fig. 2b) enable efficient proton transport [25]. In some ways, this design resembles the small pores in wood or the interstitial spaces between cells [26, 27], making clay sheets ideal for nature-inspired nanoengineering.
2.2 Clays as natural 2D materials
To identify the best clay materials for the blue device, we tested the electrical and hydrophilic properties of several abundant clays, including kaolinite, illite, and smectite (see SI Fig. S1). These clays naturally form small crystals (Fig. 3a) that can grow to micrometer sizes (SI Figs. S3-S6). Found across all continents (see SI Fig. S2) and even on Mars [28], clays consist of 1-nm-thick crystal sheets made of silica and alumina planes (Fig. 3b) [29]. The sheets are weakly bound, allowing interlayer water penetration and swelling, and can be reconstructed to form large-scale heterostructures.
To ensure the material’s purity and to exclude the dissolution of foreign species into the water channels, all the raw samples were carefully cleaned before use by multiple centrifugation and dialysis, followed by energy dispersive X-ray elemental analysis (see Methods). The water pH and conductivity were measured before and after the measurements. Finally, water was introduced to the channels by adsorption from saturated water vapor excluding foreign species presence.
The hydrophilic properties of purified clays were tested using X-ray diffraction (XRD) by assessing interlayer spacing between periodic crystal sheets. Among the studied clays, smectite exhibited the highest swelling and was selected for further experiments. The wet smectite has an interlayer distance expansion of about 1 nm compared to dry one (Fig. 3c and SI Figs. S5, S6), consistent with previous studies [30, 31]. Thus, nearly 50% of the wet clay volume consists of water-filled pores, with the remaining portion being the solid matrix.
Dielectric properties of clays were compared for dry and wet conditions by electrical conductivity spectrum measurements. Dry clays were found to be good insulators (Fig. 3d), while cleaned and hydrated samples exhibited ionic conductivity far exceeding that of bulk water (see blue dashed line for reference). Smectite, in particular, showed the highest direct-current conductivity at 0.6 S/m—comparable to polymer electrolyte membranes like Nafion [32] and bulk ionic liquids [33]. Thus, it was assigned as the best for the design of the blue device.
Enhanced proton transport in hydrated clay can be understood by considering dynamic interactions between water and the clay-crystal walls under ultraconfinement. In the crystal structure, natural substitution of Si4+ and Al3+ with Al3+ and Mg2+ [29] creates an excess negative surface charge (SI Table S1). In nature, this charge is balanced by adsorbed alkaline cations [29], but in cleaned clays, hydrogen ions (H+) from water take over this role. This increases conductivity by disrupting water’s self-dissociation equilibrium [34], raising the concentration and mobility of H3O+ and OH-. Thus, the ultraconfinement of water in clays transforms the former into a partially ionic liquid with hydronium and hydroxyl ions as charge carriers without increasing chemical complexity.
The maximum surface conductivity of ultraconfined water is expected in pores around twice the Bjerrum length nm—which marks the balance between electrostatic interactions and thermal energy (kBT). Indeed, experimentally, peak proton conductivity has been reported in water confined within nanopores of around 2 nm in various materials [12, 14, 15].
2.3 Energy storage in 1-nm-high water channels
To fabricate the blue device, we used a nanopore filtration technique (Fig. 4a). We assembled layer-by-layer graphene-clay composite electrodes separated by a pure clay membrane (Fig. 4b) from corresponding colloidal suspensions (see Methods). We produced dozens of membrane-electrode units (MEUs) with varying layer thicknesses and graphene-clay compositions (see SI Table S2). All MEUs (Fig. 4c) demonstrated excellent reproducibility, mechanical stability, and flexibility.
Fig. 4d shows a cross-sectional image of a MEU and Figs. 4e, f demonstrate corresponding EDX mapping. This analysis confirms the presence of only native clay elements (O, Al, Si) and graphene (C), as prescribed following the preliminary cleaning procedure (see SI Figs. S9, S10 for further data). Electron microscopy (SEM and STEM) shows (Fig. 4, g to i,) that reconstructed clay sheets align to form parallel channels, intersecting at domain edges and providing a pathway for molecular and ionic transport.
To test the device, the nanopores were loaded with water by exposure of MEU to the saturated water vapor. Two graphite current collectors were applied to both sides of the MEU (Fig. 5a) to avoid direct contact of metal connectors with water. The cell underwent standard electrochemical tests within a voltage window up to 2.1 V (see Methods). The charge-discharge cycles (Fig. 5b) and current-voltage characteristics (Fig. 5c) were measured, and the capacitance, coulombic efficiency, and energy efficiency were estimated. Fig. 5d shows the MEU charge-discharge mechanism by separating water’s hydronium and hydroxyl ions.
The cell showed impressive parameters and durability with no visible degradation after 60,000 cycles (Fig. 5e). This cycle life translates to over 150 years of operation counting one charge-discharge cycle per day, e.g., in the case of use with solar panels. It retains and releases energy with up to 100% Coulombic and 60% energy efficiency, due to the absence of side reactions. The initial ”annealing” period was assigned to the electric-field-stimulated wetting of nanopores.
We found that the device has an electrolysis threshold of 1.65 V (Fig. 5f, and Fig. S11), higher than the standard 1.23-V limit for bulk water at neutral conditions, which yields higher capacitance, and energy density than that in bulk water systems. This is likely due to the combined effects of high water purity and the ultraconfinement of water, which alters its electrodynamics, particularly conductivity (Fig. 3f) and polarizability [12, 13, 15]. Our system achieves a capacitance of around 40 F/g and an energy density of around 10 Wh/kg of electrode material, competitive with that of commercial supercapacitors, hybrid capacitors, and some batteries [35]. Both parameters have the potential for further optimization.
3 Discussion
We have introduced a scalable and safe platform—the ”blue device”—for electric energy storage in ultraconfined water, leveraging its unique electrodynamic properties. Unlike conventional aqueous supercapacitors, our system takes a fundamentally different approach by enabling topological crosslinking of interfacial water layers across electrodes and separators. This design, achieved using abundant materials and simple fabrication methods, offers high performance, safety, and long-term sustainability without relying on scarce resources. It exemplifies nature-inspired principles in nanoengineering.
4 Methods
4.1 Samples preparation and cleaning
Several clay types: bentonite, montmorillonite, illite, kaolinite, and silt, were purchased from Merck and Nagra in the form of powder and grains. The samples were carefully cleaned before use. To do this, the samples were preliminary milled to the sub-micron sizes. Then, clay samples were washed out through a multi-step cleaning process involving repeated dissolution, natural sedimentation preselection, washing and centrifugation cycles, dialysis, and sonication, with zero chemicals used. The resulting product was a nonprecipitating colloidal solution of clay nanoparticles in distilled water with a concentration of about 0.1 mg/ml, a bulk electrical resistivity of approximately 18 M, and a pH of 7. An electrochemically exfoliated graphene solution in distilled water was purchased from XFNano and used as received. Milli-Q water, with an 18 M resistivity, served as the water source. Polymer-filter substrates with nanopores ranging from 20 to 100 nm were purchased from Merck and used without modification.
4.2 Vacuum filtration
Clay films and devices were fabricated using vacuum filtration of colloidal graphene and pre-treated clay solutions through the nanopore polymer-filter-substrate vacuum filtration. Before use, all solutions were sonicated at 1200 Watts and 20-25 kHz and naturally thermally equilibrated with the lab temperature. For electrode fabrication, either a pure graphene solution or a clay-graphene suspension mixture was used (see SI Table S2 for different MEU compositions). A pure clay solution was utilized for the separator. To create the three-layer electrode-separator-electrode structure, the electrode solution, separator solution, and second electrode solution were sequentially filtered. A water vacuum pump facilitated water transport through the filter, and filtration proceeded step-by-step, waiting for the complete passage of water from the previous step. Conductivity and pH tests of the filtrate confirmed no contamination.
Unlike pressing dry powders, vacuum filtration produces highly dense samples without large pores, featuring unidirectionally oriented ”brick-and-mortar” structure layers of 2D materials held together by vdW forces. This method ensures remarkable mechanical stability, flexibility, and reproducibility for each membrane-electrode assembly.
4.3 Electrochemical measurements
Vacuum-filtrated samples were placed in a saturated water vapor atmosphere for 48 hours, with water saturation confirmed by mass gain and swelling checks, ensuring no contamination by foreign ionic or molecular species. Conductivity, charge-discharge (CD), cyclic voltammetry (CV), and long-term CD tests were conducted using the BioLogic SP-300 and CH Instruments devices. The voltage window was set to 2.1 V for preliminary tests and to 1.6 V for main measurements, including long-term stability tests. The second value was chosen according to the electrolysis limit observed via a hydrogen-evolution-reaction (HER) indicator in CV and CD curves (SI Fig. S13). Two graphite current collectors were applied to the electrodes in all measurements. To prevent water evaporation, Parafilm was used to seal the membrane-electrode unit. Notably, no mass change in the samples was detected over one month after the fabrication. Pressure applied to the current collectors was controlled by a custom-made sample holder and was adjusted to ensure good electrical contact.
4.4 X-ray diffraction measurement
X-ray diffraction (XRD) measurements were conducted using a Bruker D8 Discover in Bragg Brentano geometry, with a Kalpha1 - monochromated beam ( = 1.5406 ) and a Lynxeye-XE detector and at the P62 SAXS/WAXS beamline of PETRA III synchrotron facility at Deutsches Elektronen-Synchrotron (DESY). We prepared two types of samples for these measurements: clay vacuum-filtered membranes and powders. For the Bruker measurements, dry samples were prepared by vacuum annealing for 24 hours, while wet samples were prepared by exposure to saturated water vapor for 48 hours. At PETRA III, the dry samples were similarly annealed, but the wet samples were prepared by directly adding water to the powders. The crystalline interlayer distance change on water adding (swelling effect) was found in both the PETRA III and lab-based XRD measurements (SI Fig. S9).
4.5 Scanning electron microscopy
Scanning electron microscopy (SEM) observation including EDX (energy dispersive X-ray spectroscopy) measurements was performed on an SEM/FIB (focused ion beam) Zeiss CrossBeam 550 microscope equipped with Oxford Instruments Ultim Max detector. SEM images were taken with 5 kV (Figs. S21, S22) and 3 kV (Figs. S23) accelerating voltage, while EDX data measurements were done at 15 kV (Figs. 3, d to g, and Figs. S9, S10). For subsequent transmission electron microscopy observation, a lamella from a smectite membrane (Figs. S24, S25) was prepared using the SEM/FIB tool Zeiss NVision 40 using an accelerated Ga+ beam at 30 kV and 5 kV.
4.6 Scanning transmission electron microscopy
Scanning transmission electron microscopy (STEM) experiments were realized on an aberration-corrected FEI Titan Themis3 microscope operated at 300 kV. An electron probe of an 8 mrad half-convergence angle was used in association with a current of 10 pA. A dwell time of 2 s was used for STEM imaging, which limited the total electron dose deposited on the sample for acquiring atomic-resolution STEM images to 460 e-/. Annular dark-field (ADF) and integrated differential phase contrast (iDPC) STEM images were collected simultaneously, using a collection angle ranging from 17 to 104 mrad for ADF and from 4 to 15 mrad for iDPC. iDPC STEM imaging allowed us to visualize the atomic structure of the crystal composed of light elements [39], with the moderate electron dose used to avoid degradation of the crystal structure during imaging.
5 Acknowledgements
We thank Keith Stevenson, Nianduo Cai, Artem Pronin, Maria Nieves López Salas, and Anton Andreev for fruitful discussions, Sylvio Haas (DESY) for help during the synchrotron radiation-based X-ray scattering experiments, Vytautas Navikas for the help with the representation of Figure 1e, and Milad Sabzehparvar for assistance in the lab. We also acknowledge the scientific exchange and support of the Centre for Molecular Water Science, and the research initiative BlueMat: Water-Driven Materials, Hamburg (Germany). We thank the EPFL Center for Electron Microscopy (CIME) for access to electron microscopes.
Funding
Y.T. acknowledges the financial support from the Swiss National Science Foundation through the National Centre of Competence in Research Bio-Inspired Materials and Grant No. 200021-192037. A.Ra. and T.H.C. acknowledge funding from the European Research Council (grant 101020445—2D-LIQUID). S.B. and L.L. are grateful for the financial support provided by the Swiss National Science Foundation (grant No. 200021-204099, Division II). Funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in the DFG Graduate School GRK 2462 ‘Processes in natural and technical Particle-Fluid-Systems (PintPFS)’ (Project No. 390794421) is gratefully acknowledged. G.T. acknowledges support from the SNSF Eccellenza Grant 194181 and the SNSF Starting Grant 211695
Declarations
There are no conflicts of interest to declare. All data is available within this paper and Supplementary Information.
Author contribution
V.A. proposed and directed the research with the help of F.B., A.Ra., G.T., and L.L.; S.B. prepared the clay materials and characterized their optical, transport, and electrical properties together with V.A. and under the supervision of L.L.; Y.T. developed the vacuum filtration protocol and assembled the membranes and membrane-electrode units with V.A.; J.M. and V.A. conducted the electrochemical measurement under the guidance of G.T.; A.Ry. carried out the data analysis; T.-H.C. develop the protocol and processed the raw materials, controlled their purity, and measured -potentials; L.N., V.B., and Y.T. conducted electron microscopy characterization; P.S. conducted X-ray diffraction experiments; M.L. and P.H. conducted synchrotron radiation-based X-ray scattering experiments and analyzed the data; V.A. wrote the manuscript; and all authors contributed to discussions.